, In a cartesian plane (OXY) the solution of the equation e2x – 2ex + 1 = 0 represent,

A.     A point

B.     A line

C.     A circle


1
Expert's answer
2021-09-14T06:05:41-0400

The equation "e^{2x}-2e^x+1=0" is equivalent to the equation "(e^x-1)^2=0." It follows that "e^x-1=0," and hence "e^x=1." We conclude that the point "x=0" is a unique solution of the equation.


Answer: A